OpenStudy (anonymous):

help please 5x - 4y = 3 2x+ 7y = -10

OpenStudy (anonymous):

Okay, so let us isolate x. 2x = -10 - 7y x = -5 - 7/2y Now we can use x in the first equation. 5(-5-7/2y) - 4y = 3 -25 - 35/2y - 4y = 3 -25 - 43/2y = 3 -43/2y = 28 -43y = 56 y = -56/43 If this helps, please leave a medal.